# Language Models Use Lookbacks to Track Beliefs

### v2

@@ Appendix B The CausalToM Dataset
Appendix C Causal Mediation Analysis
@@
Appendix D Desiderate Based Patching Via Causal Abstraction
@@
Appendix E Pseudocode for the Belief Tracking High-Level Causal Model
@@
Appendix F Desiderata-based Component Masking
@@
Where $\lambda$ is a hyperparameter used to control the rank of the subspace and $\mathbf{m}$ is the learnable mask. See Appendix D for details on how the causal model output under intervention are computed. We trained $\mathbf{m}$ for one epoch with ADAM optimizer, on batches of size $4$ and a learning rate of $0.01$. During training, the parameters of $\mathbf{m}$ are continuous and constrained to lie within the range $[0,1]$. To enforce this constraint, we clamp their values after each gradient update. During evaluation, we binarize the mask by rounding each parameter to the nearest integer, i.e., 0 or $1$.
Appendix G Aligning Character and Object OIs
@@
Appendix H Aligning Query Character and Object OIs
@@
Appendix I Speculated Payload in Visibility Lookback
@@
Fig.18 shows the experimental results. Knocking out any of the previous sentences affects the model’s ability to produce the correct output. The decrease in IIA in the early layers can be explained by the restriction on the movement of character OIs. Specifically, the second visibility sentence mentions the first and second characters, whose character OIs must be fetched before the model can perform any further operations. Therefore, we believe the decrease in IIA until layer 15, when the character OIs are formed (based on the results from Section G), can be attributed to the model being restricted from fetching the character OIs. However, the persistently low IIA even after this layer—especially when both the second and first visibility sentences are involved—indicates that some vital information is being fetched by the second visibility sentence, which is essential for forming the coherent Visibility lookback payload. Thus, we speculate that the Visibility payload encodes information about the observed character, specifically their character OI, which is later used to fetch the correct state OI.
Appendix J Correlation Analysis of Causal Subspaces and Attention Heads
@@
Appendix K Belief Tracking Mechanism in BigToM Benchmark
@@
To justify the presence of OIs, we conduct an interchange intervention experiment, similar to the one described in Section H, aiming to localize the character OI at the character token in the question sentence. We construct an original sample by replacing its question sentence with that of a counterfactual sample, selected directly from the unaltered BigToM dataset. Consequently, when processing the original sample, the model has no information about the queried character and, as a result, produces unknown as the final output. However, if we replace the residual vector at the queried character token in the original sample with the corresponding vector from the counterfactual sample (which contains the character OI), the model’s output changes from unknown to the state token(s) associated with the queried object. This is because inserting the character OI at the queried token provides the correct pointer information, aligning with the address information at the correct state token(s), thereby enabling the model to form the appropriate QK-circuit and retrieve the state’s OI. As shown in Fig. 21, we observe a high IIA between layers 9289-28—similar to the pattern seen in CausalToM—suggesting that the queried character token encodes the character OI in its residual vector within these layers.
@@
Appendix L Generalization of Belief Tracking Mechanism on CausalToM to Llama-3.1-405B-Instruct
This section presents all the interchange intervention experiments described in the main text, conducted using the same set of counterfactual examples on Llama-3.1-405B-Instruct, using NDIF  Fiotto-Kaufman et al. (2025). Each experiment was performed on 80 samples. Due to computational constraints, subspace interchange intervention experiments were not conducted. The results indicate that Llama-3.1-405B-Instruct employs the same underlying mechanism as Llama-3-70B-Instruct to reason about belief and answer related questions. This suggests that the identified belief-tracking mechanism generalizes to other models capable of reliably performing the task.
@@
Figure 25: Payload and address of Binding lookback
@@
Figure 26: Source Information of Binding lookback
@@
Figure 27: Source Reference Information of Binding lookback without freezing address and payload
@@
Figure 28: Character OI
@@
Figure 29: Object OI
@@
Figure 30: Query Object OI
@@
Figure 31: Query Character OI
@@
Figure 32: Answer Lookback Pointer and Payload
@@
Figure 33: Visibility Lookback

### v3

@@ Appendix B The CausalToM Dataset
Appendix C Mechanistic Interpretability Concepts
In this section, we summarize the residual-stream framework on transformer architecture as well as describe QK- and OV-circuits, which form the conceptual foundation for the lookback mechanisms analyzed in this work (Elhage et al., 2021).
C.1 The Residual Stream Framework
Following the standard transformer-circuits framework (Elhage et al., 2021), we adopt the residual-stream view, in which every layer writes its contributions into a shared vector space. For a token at position $t$ and layer $\ell$, let $\mathbf{r}_{t}^{(\ell)}\in\mathbb{R}^{d}$ denote its residual-stream representation. The embedding layer initializes $\mathbf{r}_{t}^{(0)}$, and each sub-layer (self-attention and MLP) adds a vector:
$\mathbf{r}_{t}^{(\ell+1)}=\mathbf{r}_{t}^{(\ell)}+\mathrm{Attn}_{t}^{(\ell)}+\mathrm{MLP}_{t}^{(\ell)}.$
This additive structure allows information written at one token position to be recovered many layers later by downstream attention heads. Within the residual stream, distinct subspaces encode different types of information; for example, LMs allocate a specific low-rank subspace for representing the OI associated with a crucial token.
C.2 QK Circuits: Routing Attention via Pointers and Addresses
Each attention head computes queries and keys from the residual stream:
$\mathbf{q}_{t}=W_{Q}\mathbf{r}_{t}^{(\ell)},\qquad\mathbf{k}_{s}=W_{K}\mathbf{r}_{s}^{(\ell)},$
and forms attention weights through scaled dot products:
$\alpha_{t\rightarrow s}\propto\exp\left(\frac{\mathbf{q}_{t}^{\top}\mathbf{k}_{s}}{\sqrt{d_{k}}}\right).$
The QK-circuit determines where the model looks. In the context of lookback mechanisms, a pointer is a component of the residual stream at the lookback token that, after the $W_{Q}$ projection, aligns strongly with an address stored in an earlier token after its $W_{K}$ projection. A high dot product between these transformed vectors causes the model to attend to the recalled token, thereby forming a QK-circuit.
C.3 OV Circuits: Transferring Payloads Across Tokens
Once the QK-circuit selects a recalled token, the attention head retrieves the value-vector payload stored in its residual stream, thereby forming an OV-circuit:
$\mathbf{v}_{s}=W_{V}\mathbf{r}_{s}^{(\ell)},$
which is then projected through $W_{O}$:
$\mathrm{Attn}_{t}^{(\ell)}=W_{O}\!\left(\sum_{s}\alpha_{t\rightarrow s}\mathbf{v}_{s}\right).$
This OV-circuit determines what information is transferred.
The QK- and OV-circuits thus together implement the mechanism by which a lookback retrieves its payload: QK selects the source of information, and OV copies the relevant content into the current token’s residual stream.
C.4 Lookback as Pointer Dereference
A lookback consists of three components:
  1. 1.
    a source reference (e.g., an OI) created at an early token;
  2. 2.
    an address copy of this reference stored in the recalled token’s residual stream and placed alongside the payload to be retrieved later;
  3. 3.
    a pointer copy stored in the lookback token’s residual stream.
When the pointer and address align to form the QK-circuit, the head dereferences the pointer and uses the OV-circuit to bring the payload forward. This general pattern recurs across the binding lookback, answer lookback, and visibility lookback mechanisms. The residual-stream view makes this structure explicit by enabling precise localization of pointers, addresses, and payloads in the model’s internal representations.
C.5 Ordering ID Assignment
The LM assigns an Ordering ID (OI; Dai et al. (2024)) to the character, object, and state tokens. These OIs, encoded in a low-rank subspace of the internal activation, serve as a reference that indicates whether an entity is the first or second of its type independent of its token value. For example, in Fig. 3a, Bob is assigned the first character OI, while Carla receives the second. We validate the presence of OIs through multiple experiments, where intervening on tokens with identical token values but different OIs alters the model’s internal computation, leading to systematic changes in the final output predicted by our high-level causal model. The LM then uses these OIs as building blocks, feeding them into lookback mechanisms to track and retrieve beliefs.
Appendix D Model Behavioral Evaluation
Table 1: Model performance under the visibility and no-visibility conditions of CausalToM (mean ± standard deviation), computed over 10 independent runs of 100 samples each. Highlighted models achieve accuracy above 80% in both conditions.
Model Name No Visibility (mean ± std) Visibility (mean ± std)
7B Models
Llama-2-7b-hf 0.011 $\pm$ 0.007 0.006 $\pm$ 0.008
Qwen2.5-7B 0.280 $\pm$ 0.039 0.061 $\pm$ 0.017
Qwen2.5-7B-Instruct 0.948 $\pm$ 0.022 0.719 $\pm$ 0.031
8B Models
Llama-3.1-8B 0.446 $\pm$ 0.046 0.293 $\pm$ 0.042
Llama-3.1-8B-Instruct 0.722 $\pm$ 0.044 0.310 $\pm$ 0.035
Meta-Llama-3-8B 0.297 $\pm$ 0.027 0.117 $\pm$ 0.024
Meta-Llama-3-8B-Instruct 0.349 $\pm$ 0.042 0.085 $\pm$ 0.025
13B Models
Llama-2-13b-hf 0.328 $\pm$ 0.041 0.117 $\pm$ 0.027
OLMo-2-1124-13B-Instruct 0.522 $\pm$ 0.034 0.360 $\pm$ 0.049
14B Models
Qwen2.5-14B 0.865 $\pm$ 0.038 0.433 $\pm$ 0.040
[…]
27B Models
Gemma-3-27b-it 0.527 $\pm$ 0.036 0.388 $\pm$ 0.034
32B Models
OLMo-2-0325-32B-Instruct 0.814 $\pm$ 0.033 0.679 $\pm$ 0.029
70B Models
[…]
405B Models
[…]
The section reports the behavioral performance of several LM families on the CausalToM task, over $10$ independent runs of $100$ samples each. As shown in Table 1, most models, particularly the smaller ones, struggle to perform the task consistently. In contrast, the Llama models, especially Llama-3-70B-Instruct and Llama-3.1-405B-Instruct, and the Qwen model Qwen2.5-14B-Instruct achieve over $80\%$ accuracy in both the visibility and no-visibility conditions. Motivated by this, we conduct a mechanistic analysis of these high-performing models to investigate how LMs represent and track beliefs.
Appendix E Causal Mediation Analysis
@@
Appendix F Desiderate Based Patching Via Causal Abstraction
@@
Appendix G Pseudocode for the Belief Tracking High-Level Causal Model
@@
Appendix H Desiderata-based Component Masking
@@
Where $\lambda$ is a hyperparameter used to control the rank of the subspace and $\mathbf{m}$ is the learnable mask. See Appendix F for details on how the causal model output under intervention are computed. We trained $\mathbf{m}$ for one epoch with ADAM optimizer, on batches of size $4$ and a learning rate of $0.01$. During training, the parameters of $\mathbf{m}$ are continuous and constrained to lie within the range $[0,1]$. To enforce this constraint, we clamp their values after each gradient update. During evaluation, we binarize the mask by rounding each parameter to the nearest integer, i.e., 0 or $1$.
Appendix I Aligning Character and Object OIs
@@
Appendix J Aligning Query Character and Object OIs
@@
Appendix K Speculated Payload in Visibility Lookback
@@
Fig.18 shows the experimental results. Knocking out any of the previous sentences affects the model’s ability to produce the correct output. The decrease in IIA in the early layers can be explained by the restriction on the movement of character OIs. Specifically, the second visibility sentence mentions the first and second characters, whose character OIs must be fetched before the model can perform any further operations. Therefore, we believe the decrease in IIA until layer 15, when the character OIs are formed (based on the results from Section I), can be attributed to the model being restricted from fetching the character OIs. However, the persistently low IIA even after this layer—especially when both the second and first visibility sentences are involved—indicates that some vital information is being fetched by the second visibility sentence, which is essential for forming the coherent Visibility lookback payload. Thus, we speculate that the Visibility payload encodes information about the observed character, specifically their character OI, which is later used to fetch the correct state OI.
Appendix L Correlation Analysis of Causal Subspaces and Attention Heads
@@
Appendix M Belief Tracking Mechanism in BigToM Benchmark
@@
To justify the presence of OIs, we conduct an interchange intervention experiment, similar to the one described in Section J, aiming to localize the character OI at the character token in the question sentence. We construct an original sample by replacing its question sentence with that of a counterfactual sample, selected directly from the unaltered BigToM dataset. Consequently, when processing the original sample, the model has no information about the queried character and, as a result, produces unknown as the final output. However, if we replace the residual vector at the queried character token in the original sample with the corresponding vector from the counterfactual sample (which contains the character OI), the model’s output changes from unknown to the state token(s) associated with the queried object. This is because inserting the character OI at the queried token provides the correct pointer information, aligning with the address information at the correct state token(s), thereby enabling the model to form the appropriate QK-circuit and retrieve the state’s OI. As shown in Fig. 21, we observe a high IIA between layers 9289-28—similar to the pattern seen in CausalToM—suggesting that the queried character token encodes the character OI in its residual vector within these layers.
@@
Appendix N Generalization of Belief Tracking Mechanism on CausalToM to Llama-3.1-405B-Instruct and Qwen2.5-14B-Instruct
This section presents all the interchange intervention experiments described in the main text, conducted using the same set of counterfactual examples on Llama-3.1-405B-Instruct and Qwen2.5-14B-Instruct, using NDIF  Fiotto-Kaufman et al. (2025). Each experiment was performed on 80 samples. Due to computational constraints, subspace interchange intervention experiments were not conducted. As illustrated in Figures 2642, the results indicate that Llama-3.1-405B-Instruct and Qwen2.5-14B-Instruct employ the same underlying mechanism as Llama-3-70B-Instruct to reason about belief and answer related questions. This suggests that the identified belief-tracking mechanism generalizes to not only larger but models in other models families as well.
@@
Figure 25: Payload and address of Binding lookback in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: coffee
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: beer
Refer to caption CounterfactualOriginal
Figure 26: Payload and address of Binding lookback in Qwen2.5-14B-Instruct.
@@
Figure 27: Source Information of Binding lookback in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: tea
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: beer
Refer to caption CounterfactualOriginal
Figure 28: Source Information of Binding lookback in Qwen2.5-14B-Instruct.
@@
Figure 29: Source Reference Information of Binding lookback without freezing address and payload in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: coffee
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: beer
Refer to caption CounterfactualOriginal
Figure 30: Source Reference Information of Binding lookback without freezing address and payload in Qwen2.5-14B-Instruct.
@@
Figure 31: Character OI in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: coffee
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: water
Refer to caption CounterfactualOriginal
Figure 32: Character OI in Qwen2.5-14B-Instruct.
@@
Figure 33: Object OI in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: coffee
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: tea
Refer to caption CounterfactualOriginal
Figure 34: Object OI in Qwen2.5-14B-Instruct.
@@
Figure 35: Query Object OI in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: coffee
Question: What does Bob believe the cup contains?
Answer: unknown
Causal Model Output: water
Refer to caption CounterfactualOriginal
Figure 36: Query Object OI in Qwen2.5-14B-Instruct.
@@
Figure 37: Query Character OI in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: coffee
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: water
Refer to caption CounterfactualOriginal
Figure 38: Query Character OI in Qwen2.5-14B-Instruct.
@@
Figure 39: Answer Lookback Pointer and Payload in Llama-3.1-405B-Instruct.
Question: What does Carla believe the cup contains?
Answer: tea
Question: What does Carla believe the cup contains?
Answer: coffee
Refer to caption CounterfactualOriginal
Figure 40: Answer Lookback Pointer and Payload in Qwen2.5-14B-Instruct.
@@
Figure 41: Visibility Lookback in Llama-3.1-405B-Instruct.
Question: What does Carla believe the bottle contains?
Answer: water
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: coffee
Refer to caption CounterfactualOriginal
Figure 42: Visibility Lookback in Qwen2.5-14B-Instruct.
Appendix O Generalization of Belief Tracking Mechanism on CausalToM with Three Character-Object-State Triples
Experimental results from the main text and the previous section show that the identified belief-tracking mechanism generalizes across both model scale and model family. Although the interchange-intervention experiments on BigToM demonstrate that this mechanism transfers to real-world benchmarks, it remains unclear whether it would still generalize when the prompt includes a larger number of characters, objects, and states.
To investigate this, we conducted behavioral evaluations of larger Llama models and Qwen models on tasks involving three characters, three objects, and three states, under both no-visibility and visibility conditions. All three models succeeded in the no-visibility condition, but none consistently solved the visibility condition. Consequently, we performed interchange-intervention experiments only for the no-visibility condition on Qwen2.5-14B-Instruct to assess whether the mechanism generalizes. As shown in Figures 4350, Qwen2.5-14B-Instruct utilizes the same mechanism involving binding and answer lookbacks to solve the no-visibility belief tracking task.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the cup contains?
Answer: tea
Causal Model Output: beer
Refer to caption CounterfactualOriginal
Figure 43: Payload and address of Binding lookback with three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: tea
Refer to caption CounterfactualOriginal
Figure 44: Source Information of Binding lookback in three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the cup contains?
Answer: coffee
Causal Model Output: tea
Refer to caption CounterfactualOriginal
Figure 45: Source Information of Binding lookback without freezing address and payload in three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the tun contains?
Answer: unknown
Causal Model Output: tea
Refer to caption CounterfactualOriginal
Figure 46: Character OI in three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the bottle contains?
Answer: unknown
Causal Model Output: coffee
Refer to caption CounterfactualOriginal
Figure 47: Object OI in three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the tun contains?
Answer: unknown
Causal Model Output: tea
Refer to caption CounterfactualOriginal
Figure 48: Query Character OI in three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with tea. Then Bob grabs an opaque bottle and fills it with beer. Finally, Carla grabs another opaque cup and fills it with coffee.
Question: What does Carla believe the cup contains?
Answer: coffee
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Tom believe the cup contains?
Answer: unknown
Causal Model Output: tea
Refer to caption CounterfactualOriginal
Figure 49: Query Object OI in three character-object-state triples in Qwen2.5-14B-Instruct.
Tom, Bob, and Carla are working in a busy restaurant. To complete an order, Tom grabs another opaque tun and fills it with wine. Then Bob grabs an opaque bottle and fills it with juice. Finally, Carla grabs another opaque cup and fills it with milk.
Question: What does Carla believe the cup contains?
Answer: milk
Bob, Carla, and Tom are working in a busy restaurant. To complete an order, Bob grabs an opaque bottle and fills it with beer. Then Carla grabs another opaque cup and fills it with coffee. Finally, Tom grabs another opaque tun and fills it with tea.
Question: What does Carla believe the cup contains?
Answer: coffee
Intervention: Answer Pointer (Refer to caption),  Causal Model Output: tea Intervention: Answer Payload (Refer to caption),  Causal Model Output: milk
Refer to caption CounterfactualOriginal
Figure 50: Answer Lookback Pointer and Payload in three character-object-state triples in Qwen2.5-14B-Instruct.